What is an equation of the line perpendicular of y= -x-2 & through (-2,4)?
Answer by nerdybill(7384)

What is an equation of the line perpendicular of y= -x-2 & through (-2,4)?
Since the slope of:
y= -x-2
is -1 a line that is perpendicular must have a slope of:
1 (negative reciprocal)
Use slope (1) and (-2,4) and plug into the "point-slope" form:
y - y1 = x(x - x1)
y - 4 = 1(x - (-2))
y - 4 = (x + 2)
y = x + 2 +4
y = x + 6